solve by elimination method and give an ordered pair.
x+2y=11
-x+5y=17

Answer by rapaljer(4671)

x+2y =11
-x+5y=17

Just add these equations together, and the x terms subtract out. Now you can solve for y:

7y=28
y=4

Now substitute this value of y=4 back into either equation. The first equation will do:

x+2y=11
x+2(4)=11
x+8=11
x=3

Check by substituting both numbers into the second equation:
-x+5y=17
-3+5(4)=17
-3+20=17
IT CHECKS!!

R^2